Dianne invests P3500, part at 3%, part at 4%. How much does she invest at each rate if she receive in one year interest of P115?
Answer by mananth(16946)

Investment I 3.00% per annum
Investment II 4.00% per annum

x+y=3500 ------------------------1
3.00%x+4.00%y=$115.00
Multiply by 100
3x+4y=$11,500.00--------2

Multiply (1) by -3
we get

-3x-3y =-10500.00

Add this to (2)

0 x +1 y = $1,000.00
y= P1,000.00 investment at 4.00%
Balance P2,500.00 investment at 3.00%